# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v1

@@ Abstract
This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a novel neural-guided meta-heuristic algorithm for combinatorial optimization. GFACS integrates generative flow networks (GFlowNets) with the ant colony optimization (ACO) methodology. GFlowNets, a generative model that learns a constructive policy in combinatorial spaces, enhance ACO by providing an informed prior distribution of decision variables conditioned on input graph instances. Furthermore, we introduce a novel combination of training tricks, including search-guided local exploration, energy normalization, and energy shaping to improve GFACS. Our experimental results demonstrate that GFACS outperforms baseline ACO algorithms in seven CO tasks and is competitive with problem-specific heuristics for vehicle routing problems. The source code is available at https://github.com/ai4co/gfacs.
GFlowNets, Ant Colony Optimization, Neural Combinatorial Optimization
@@ 1 Introduction
Combinatorial Optimization Problems (COPs) constitute a category of optimization problems focused on finding the optimal arrangement of discrete structures, such as sets, sequences, and graphs, that maximizes a given objective function. COPs are prevalent in various fields, including compiler scheduling (Zhang et al., 2022b) and hardware design (Kim et al., 2023b), but they are usually NP-hard. Over the decades, heuristics that quickly find near-optimal solutions have been developed to tackle this NP-hardness. However, these problem-specific heuristics require considerable effort and domain knowledge to make them effective. Thus, meta-heuristic algorithms, such as Ant Colony Optimization (ACO) (Dorigo et al., 2006), have been proposed to provide a more generalized approach to designing heuristics.
ACO is a meta-heuristic algorithm inspired by the foraging behaviors of ants, and it has demonstrated effectiveness in solving COPs and general optimization problems. In ACO, a population of artificial ants iteratively builds solutions and updates ‘pheromone trails’ to explore the solution space. At each iteration, solutions are constructed through stochastic moves guided by instance-specific pheromone trails and problem-specific heuristic measures indicating the goodness of sampling certain trails; see Figure 1 for an illustration. While the pheromone trails are updated based on solutions constructed during the iterative procedure of solving an instance, the heuristic measures typically need to be predefined in a matrix form (e.g., heat map). However, developing an effective heuristic matrix for ACO involves expert knowledge and manual tuning, posing challenges in less-explored problem domains where the expertise is limited.
Figure 1: Illustration of ACO. Edge probability is obtained from the multiplication of heuristics and pheromone trails. Ants constructively complete solutions based on the edge probability, then update the pheromone trail with the constructed samples.
Recently, a neural-guided Ant Colony Optimization (ACO) method, DeepACO, was introduced by Ye et al. (2023). This technique utilizes graph neural networks (GNN) (Joshi et al., 2021) to model the heuristic matrix of input instance graphs of COPs. Specifically, this framework is developed as a GNN-based policy that selects an edge (action) based on previously chosen edges (state). Then, the composited policy generates a state trajectory from an empty set (initial state) to a complete solution (terminal state), as in Figure 1. DeepACO incorporates reinforcement learning (RL) to train the GNN-based policy model with a reward signal from the terminal state of the generated trajectory. After RL training, the learned GNN is utilized to construct the heuristic matrix for a given problem instance, thereby providing a good prior to ACO to explore the solution space more efficiently. Note that the trained GNN acts as a parameterized prior, effectively transferring the solution exploration strategy obtained during training to each new instance.
However, we observe certain limitations in DeepACO, primarily in three aspects. The first limitation ( ) is that the usage of a behavior policy is limited since DeepACO uses on-policy RL to train the graph representation. It potentially weakens the performance of heuristic matrices by limiting the choice of behavior policy, which is crucial for the effective exploration of massive combinatorial spaces. Secondly ( ), the reward-maximizing RL of DeepACO fails to satisfactorily capture the diversity of solutions, i.e., the multi-modality of the solution landscape. In CO, a diversity of good candidates is highly advantageous, as it increases the chances of finding a global optimum after a local search; see Figure 2. Lastly ( ), the inherent symmetry in combinatorial solutions, where multiple trajectories can lead to identical combinatorial outcomes, is challenging to consider in the context of RL. For example, in the Traveling Salesman Problem (TSP), distinct paths like (1,2,3,4) and (4,3,2,1) represent the same cyclic solution. As DRL methods train policy on a tree-structured MDP, each trajectory is identified as a different solution.
Figure 2: Importance of capturing multi-modality in combinatorial optimization with local search post-processing ( ). The probability of improved solutions rises after local search, even from initially higher-cost samples, particularly with a sampler adept at capturing multi-modality.
Contribution.
@@
This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets (Bengio et al., 2021, 2023). First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in COPs. Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms and well-suited for modeling the multi-modality nature of certain distributions. Lastly ( ), GFlowNets, which model the solution generative process on a directed acyclic graph (DAG), incorporate the symmetricity of combinatorial solutions by design (Bengio et al., 2023).
To adequately integrate GFlowNet into ACO, we propose an effective combination of training techniques. In our approach, we employ a strategic guided exploration with two distinct behavior policies to collect exploratory and exploitation samples. One aims for a diverse range of samples, and the other aims for locally optimized samples. This approach effectively addresses the over-exploration issue in GFlowNets, which often results in poor performances due to an excessive focus on diversity. To encourage further exploration, we introduce an energy reshaping to compensate samples that have initially high energy but can be significantly improved, as exemplified by the red samples in Figure 2. At last, a shared energy normalization is employed to stabilize the training of conditional GFlowNets, wherein the energy landscape is contingent upon the variability of input problems.
@@
Our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks, which vehicle routing problems, scheduling problems, and a grouping problem. Moreover, GFACS outperforms or shows similar performance to competitive deep reinforcement learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$.
@@ 2.1 Ant Colony Optimization
Ant Colony Optimization (ACO) is a meta-heuristic inspired by ant behavior in finding optimal paths. Traditional ACO algorithms, such as Ant System (AS) (Dorigo et al., 1996) and MAX-MIN AS (Stützle & Hoos, 2000), rely on specific heuristics for success. To reduce the reliance on heuristic design, Ye et al. (2023) introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
@@ 2.2 Neural Combinatorial Optimization
Neural combinatorial optimization (Bello et al., 2016) is a deep learning-based method to train an instance-conditional policy using supervised learning (Vinyals et al., 2015; Fu et al., 2021; Sun & Yang, 2023) or unsupervised learning (Khalil et al., 2017; Kool et al., 2018; Kwon et al., 2020). Usually, supervised learning methods are beneficial when the target problem has an accurate but time-consuming expert solver. The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize them, allowing for scalability and faster inferences than the expert solver. On the other hand, unsupervised learning proceeds through self-guided trial and error, using the COP objective as a training signal. This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable (Bello et al., 2016; Ahn et al., 2020). Our method can be categorized as an unsupervised learning method, which does not need expert labels. See Appendix F for a more comprehensive survey.
@@ 2.3 GFlowNets and Combinatorial Optimization
GFlowNets, tailored for sampling combinatorial objects proportionally to a given reward function, play a pivotal role in diversifying the candidate pool for molecule generation in crucial areas like drug discovery (Bengio et al., 2021). Recent efforts to advance GFlowNets have focused on developing new training objectives (Malkin et al., 2022; Madan et al., 2022; Zhang et al., 2023b), enhancing credit assignment (Pan et al., 2023a; Jang et al., 2024), and improving exploration strategies (Pan et al., 2022; Rector-Brooks et al., 2023; Pan et al., 2023a; Kim et al., 2023d, c). Our contribution lies in proposing an efficient exploration technique tailored for combinatorial optimization problems.
GFlowNets have also found successful applications in some COPs (Zhang et al., 2022b, 2023a). Zhang et al. (2023a) effectively addresses the symmetric nature of graph constraint optimization, showcasing remarkable performance in graph CO tasks like maximum independent sets. Zhang et al. (2022b) solves proxy-based scheduling problems. Despite their success, their method has limited applications to either locally decomposable graph CO problems (Ahn et al., 2020) or robust scheduling problems. Our method is a neural-guided meta-heuristic, which aims to be a general-purpose CO solver. Additionally, this paper appears to be the first to address the traveling salesman problem and its variations using GFlowNets. These tasks present a significant challenge due to the global constraints to compare with graph CO tasks.
@@ 3 Preliminary
3.1 Travelling Salesman Problem
The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem (CVRP) and orienteering problem (OP). In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example. Note that the formulations can be extended to other COPs straightforwardly.
A TSP instance is defined on a (fully connected) graph $\mathcal{G}=\{\mathcal{V},\mathcal{D}\}$, where $\mathcal{V}$ and $\mathcal{D}$ are a node set and edge set, respectively. Each node $i\in\mathcal{V}$ has two-dimensional coordinates $v_{i}\in[0,1]^{2}$ as a feature, and each edge $(i,j)\in\mathcal{D}$ has a feature $d_{i,j}$ which is the distance between node $i$ and $j$, i.e., $d_{i,j}=\|v_{i}-v_{j}\|_{2}$. A solution to TSP, $x(\mathcal{G})$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$. The tour length is computed as $f(x(\mathcal{G}))=\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}+d_{\pi_{N},\pi_{1}}$. The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $f(\cdot,\mathcal{G})$.
@@ 3.2 Ant Colony Optimization
At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Equation 1. The pheromone trails $\rho$ are then adjusted using the solution set and corresponding costs, thereby updating the solution construction probability. The pheromone update rules vary depending on the ACO algorithm used. For example, in Ant System (AS) (Dorigo et al., 1996), all $K$ solutions contribute to the pheromone update:
@@
Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process, only serving as a prior. Thus, ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al. (2006) for details. In DeepACO (Ye et al., 2023), the learned GNN, parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for exploring the solution space.
Figure 3: Solution sampling process of GFACS. The GNN is trained with GFlowNets and serves as an expert heuristic in GFACS.
@@ 3.3 Generative Flow Networks
Generative Flow Networks, or GFlowNets (Bengio et al., 2021, 2023) provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$. GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object. All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
We define a complete trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$. We denote $\tau_{{}_{\rightarrow}x}$ when the terminal state $s_{n}$ is finalized to the corresponding object $x\in\mathcal{X}$. We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$. Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$. We can use the forward policy to compute in a forward way the probability of a trajectory:
$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$ (3)
Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x$:
$P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$ (4)
Trajectory balance (Malkin et al., 2022).
Trajectory balance is the most widely used learning objective in a GFlowNet framework. When the trajectory balance objective is minimized, we achieve that the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{\tau_{\rightarrow x}\in\mathcal{T}}P_{F}(\tau)$. The trajectory balance loss $\mathcal{L}_{\text{TB}}$ works by training three models, a learnable scalar $Z_{\theta}$ for the partition function, a forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau_{\rightarrow x}|x;\theta)$ to minimize the following objective:
$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}$ (5)
4 Generative Flow Ant Colony Sampler

### v2

@@ Abstract
This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a neural-guided probabilistic search algorithm for solving combinatorial optimization (CO). GFACS integrates generative flow networks (GFlowNets), an emerging amortized inference method, with ant colony optimization (ACO), a promising probabilistic search algorithm. Specifically, we use GFlowNets to learn a constructive policy in combinatorial spaces for enhancing ACO by providing an informed prior distribution over decision variables conditioned on input graph instances. Furthermore, we introduce a novel off-policy training algorithm for scaling conditional GFlowNets into large-scale combinatorial spaces by leveraging local search and shared energy normalization. Our experimental results demonstrate that GFACS outperforms baseline ACO algorithms in seven CO tasks and is competitive with problem-specific heuristics for vehicle routing problems.
@@ 1 Introduction
Combinatorial optimization (CO) is a crucial domain involving high-impact applications, including vehicle routing, hardware design optimization, and biological sequence design. Ant colony optimization (ACO) [14, 68] is a general-purpose meta-heuristic algorithm designed to address these tasks. ACO employs a constructive solution generation scheme that sequentially adds elements to a partial solution, known for its flexibility across CO tasks. This approach involves collecting experiences from multiple agents (ants) to iteratively update a distribution (pheromone update) based on a prior distribution known as the heuristic matrix. The major benefits of ACO include: (1) its robustness against local optima, enabled by the algorithm’s stochastic nature, which dynamically balances exploration and exploitation, preventing premature convergence [14]; and (2) its flexible solution generation scheme, which allows it to effectively solve a wide range of combinatorial optimization tasks [79].
The main challenge with ACO lies in designing effective prior distributions (i.e., the heuristic matrix) for their probability models. Inadequate priors can lead to time-consuming pheromone update iterations. Recent advancements, known as DeepACO [79], address this by using policy gradient reinforcement learning (RL) [76, 71] to pretrain the priors conditioned on given problem instances. RL trains these networks to guide highly rewarded regions, allowing generalization to new problems and enhanced performance.
Motivated by DeepACO, we propose a deep learning-based prior distribution design. However, this paper tackles challenges related to the design of prior distribution through deep amortized inference rather than the policy gradient RL method. This is motivated by the following three primary limitations in DeepACO’s use of policy gradient RL. First ( ), the on-policy nature of DeepACO restricts the flexible use of a behavior policy, potentially undermining performance by limiting effective exploration of vast combinatorial spaces. Second ( ), the reward maximizing form of RL in DeepACO struggles to capture the diverse solutions necessary in CO, even though having a variety of good candidates could significantly enhance the likelihood of finding a global optimum after local searches. Last ( ), the inherent symmetry in combinatorial solutions—where different paths can lead to identical outcomes—presents a significant challenge not adequately addressed in DeepACO.
Contribution. This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets [4, 5]. Our twofold contribution can be summarized as follows. (1) We introduce a new viewpoint to ACO using amortized inference instead of policy gradient RL. (2) Additionally, we present a novel GFlowNet training algorithm for large scale combinatorial tasks, applicable to a range of further uses over ACO.
@@
  • ACO as off-policy amortized inference with GFlowNets. There are three reasons that GFlowNets can improve over the existing DeepACO algorithm. First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in CO problems. Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms that are well-suited for modeling the multimodal nature of certain distributions. Lastly ( ), GFlowNets model solutions as a generative process on a directed acyclic graph (DAG), which can represent multiple trajectories that share identical states. This approach allows for accurate probabilistic modeling by considering the symmetry of combinatorial solutions [5].
  • Scaling conditional GFlowNets into large-scale combinatorial space. GFlowNets often struggle with limited scalability due to challenging exploration problems [58, 39], credit assignment issues [59, 26], and unstable training when inferring conditional distributions [82]. In combinatorial optimization, both limitations must be addressed because we deal with (1) large-scale combinatorial spaces and (2) conditional inference on problem instances. This study introduces a new off-policy training method for GFlowNets in large combinatorial spaces by effectively leveraging local search operators. It also proposes a novel training method to address instability in training conditional GFlowNets by properly normalizing conditional reward values with sample mean.
@@
In comparative experiments, our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks – vehicle routing problems, scheduling problems, and a grouping problem. Moreover, GFACS outperforms or shows similar performance to competitive unsupervised learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$. Lastly, we compared our approach with previous GFlowNet methods for solving combinatorial optimization problems [82, 84] in vehicle routing tasks, which are known to pose challenges for credit assignment due to their locally undecomposable properties, as highlighted by Ahn et al. [1].
@@ 2.1 Ant colony optimization
Ant Colony Optimization (ACO) is a meta-heuristic inspired by ant behavior in finding optimal paths. Traditional ACO algorithms, such as Ant System (AS) [14] and MAX-MIN AS [68], rely on specific heuristics for success. To reduce the reliance on heuristic design, Ye et al. [79] introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
@@ 2.2 Neural combinatorial optimization
Neural combinatorial optimization [3] is a deep learning-based method to train an instance-conditional policy using supervised learning [75, 20, 70] or unsupervised learning [30, 41, 43, 54]. Usually, supervised learning methods are beneficial when the target problem has an accurate but time-consuming expert solver. The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize the search process, converting an expensive run-time optimization into an also expensive but one-time cost of training the neural net, allowing for scalability and faster run-time inferences than the expert solver. On the other hand, unsupervised learning proceeds through self-guided trial and error, using the COP objective as a training signal. This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable [3, 1]. Our method can be categorized as an unsupervised learning method, which does not need expert labels. See Appendix F for a more comprehensive survey.
@@ 2.3 GFlowNets and combinatorial optimization
GFlowNets are designed for sampling combinatorial objects proportionally to a given reward function [4]. Theoretical studies have connected GFlowNets with variational inference [53] and maximum entropy reinforcement learning [56, 13]. Recent algorithmic advancements in GFlowNets have focused on developing new training objectives [52, 51, 85], enhancing credit assignment [59, 26], and improving exploration strategies [58, 63, 59, 39, 38].
GFlowNets have also found successful applications in some combinatorial optimization problems (COPs) [82, 84]. Zhang et al. [84] effectively address the symmetric nature of graph constraint optimization, demonstrating remarkable performance in tasks such as finding maximum independent sets. Despite their success, there are limitations in expanding to other combinatorial tasks due to their reliance on local credits for the learning objective proposed by Pan et al. [59]. Conversely, Zhang et al. [82] address proxy-based scheduling problems by developing a problem-specific neural architecture and introducing novel training methods for general-purpose conditional GFlowNets. One of their significant contributions is the proposal of a loss function based on the sample variance of the log partition function, known as VarGrad, which enhances the training process and performance.
@@ 3 Preliminary
3.1 Traveling salesman problem
The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem (CVRP) and orienteering problem (OP). In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example.
A TSP instance is defined on a (fully connected) graph $\mathcal{G}=\{\mathcal{V},\mathcal{D}\}$, where $\mathcal{V}$ and $\mathcal{D}$ are a node set and edge set, respectively. Each node $i\in\mathcal{V}$ has two-dimensional coordinates $v_{i}\in[0,1]^{2}$ as a feature, and each edge $(i,j)\in\mathcal{D}$ has a feature $d_{i,j}$ which is the distance between node $i$ and $j$, i.e., $d_{i,j}=\|v_{i}-v_{j}\|_{2}$. A solution to TSP, $x(\mathcal{G})$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$. The tour length is computed as $\mathcal{E}(x(\mathcal{G}))=d_{\pi_{N},\pi_{1}}+\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}$. The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $\mathcal{E}(x(\mathcal{G}))$.
@@ 3.2 Ant colony optimization
At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Equation 1. The pheromone trails $\rho$ are then adjusted using the solution set and corresponding costs, thereby updating the solution construction probability. The pheromone update rules vary depending on the ACO algorithm used. For example, in Ant System (AS) [14], all $K$ solutions contribute to the pheromone update:
@@
$\Delta^{k}_{i,j}=\begin{cases}C/f(\pi^{k})&\text{if $(i,j)\in\pi^{k}$}\\ 0&\text{otherwise}\end{cases}$ (3)
where $C$ is a constant. The smaller the cost $f(\pi^{k})$, the more pheromone is laid on the edges of $\pi^{k}$.
Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process; it only serves as a prior. ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al. [15] for details. In DeepACO [79], the learned graph neural network (GNN), parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for guiding probabilistic solution search process of ACO.
Figure 1: Solution sampling process of GFACS. The GNN, pretraind with a GFlowNet loss, serves as an expert heuristic to pick the next action to construct a solution, such as a tour in the TSP.
@@ 3.3 Generative flow networks
Generative Flow Networks, or GFlowNets [4, 5] provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$ towards the objective of generating $x$ with probability proportional to a given non-negative reward function $R(x)$. GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object. All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
We define a complete trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$. We denote $\tau_{{}_{\rightarrow}x}$ when the terminal state $s_{n}$ is finalized to the corresponding object $x\in\mathcal{X}$. We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$. Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$. We can use the forward policy to compute in a forward way the probability of a trajectory:$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$.
Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x=s_{n}$: $P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$.
Trajectory balance [52]. Trajectory balance is the most widely used training objective for GFlowNets. When the trajectory balance objective is minimized, we achieve that the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{\tau_{\rightarrow x}\in\mathcal{T}}P_{F}(\tau)$. The trajectory balance loss $\mathcal{L}_{\text{TB}}$ works by training three models, a learnable scalar $Z_{\theta}$ for the partition function, a forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau_{\rightarrow x}|x;\theta)$ to minimize the following objective:
$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}.$ (4)
4 Generative flow ant colony sampler (GFACS)
In this section, we provide the inference procedure for the Generative Flow Ant Colony Sampler (GFACS). GFACS is a sampling algorithm that enhances Ant Colony Optimization (ACO) by integrating it with GFlowNets. The central concept is to use GFlowNets to generate an appropriate instance-conditioned prior (edge) distribution (heuristic matrix), which more effectively guides the probabilistic search procedure of ACO.